A baker wants to make a 15 pound blend of flour that costs $.60 per pound. The blend is made by using rye flour that costs $.70 per pound and wheat flour that costs $.40 per pound. How many pounds of each flower should be used?
Answer by stanbon(75887)

6): A baker wants to make a 15 pound blend of flour that costs $.60 per pound. The blend is made by using rye flour that costs $.70 per pound and wheat flour that costs $.40 per pound. How many pounds of each flower should be used?
---------
Quantity Eq:: r + w = 15 lb
Value Eq:: 70r+40w = 60*15
---------
4r + 4w = 4*15
7r + 4w = 6*15
------
Subtract and solve for "r"::
3r = 2*15
r = 10 lbs (amt. of rye needed)
w = 15-10 = 5 lbs (amt. of wheat needed)
----------------------
Cheers,
Stan H.
-------------